A truck heading east has an initial velocity of 6 m/s. It accelerates at 2 m/s2 for 12 seconds. What distance does the truck travel in the given time?
1
Expert's answer
2020-12-10T11:06:35-0500
"s=vt+0.5at^2\\s=(6)(12)+0.5(2)(12)^2=216\ m"